# Linear Bandits with Partially Observable Features

### v2

@@ Abstract
We study the linear bandit problem that accounts for partially observable features. Without proper handling, unobserved features can lead to linear regret in the decision horizon $T$, as their influence on rewards is unknown. To tackle this challenge, we propose a novel theoretical framework and an algorithm with sublinear regret guarantees. The core of our algorithm consists of: (i) feature augmentation, by appending basis vectors that are orthogonal to the row space of the observed features; and (ii) the introduction of a doubly robust estimator. Our approach achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$, where $d$ denotes the dimension of the observed features, and $d_{h}$ represents the number of nonzero coefficients in the parameter associated with the reward component projected onto the subspace orthogonal to the row space spanned by the observed features. Notably, our algorithm requires no prior knowledge of the unobserved feature space, which may expand as more features become hidden. Numerical experiments confirm that our algorithm outperforms both non-contextual multi-armed bandits and linear bandit algorithms depending solely on observed features.
@@ 1 Introduction
We consider a linear bandit problem where the learning agent has access to only a subset of the features, while the reward is determined using the complete set of features, including both observed and unobserved elements. Conventional linear bandit problems rely on the assumption that the rewards are linear to only observed features, without accounting for the potential presence of unobserved features. However, in many real-world applications, rewards are often affected by the unobserved—hence latent—features that are not observable to the agent. For example, in online advertising without personalization—i.e., when serving general users—each advertisement is associated not only with observable features such as its category, format, or display time, but also with unobservable factors such as emotional appeal, creative design quality, and brand familiarity. These latent factors influence users’ click-through rates, yet they are not directly quantifiable. Similarly, in clinical trials, treatment outcomes depend not only on observable features like dosage and formulation, but also on unobserved factors such as manufacturing variability and potential side effect risks. Hence, accurately incorporating latent features is essential for providing precise recommendations.
@@
To address latent features,  Park & Faradonbeh (2022, 2024); Kim et al. (2023a); Zeng et al. (2025) typically assume that the true features follow a specific distribution, such as a Gaussian distribution. Establishing a regret bound sublinear in the decision horizon without such structural assumptions on the latent features remains a significant challenge and has not been accomplished yet. Key challenges in the bandit problem with partially observable features arise from the complete lack of information on the latent features. Indeed, we do not even know whether an agent observes features partially nor whether we should use the latent features. We tackle these challenges by proposing a novel linear bandit algorithm that is agnostic to partially observability. Despite having no knowledge of the unobserved features, our algorithm achieves a tighter regret bound than both linear bandit algorithms that consider only observed features and multi-armed bandit (MAB) algorithms that ignore features entirely. For brevity, we will refer to linear bandit algorithms relying solely on observed features as “OLB algorithms” henceforth. Specifically, our proposed algorithm achieves a $\sqrt{T}$-rate regret bound, without requiring any prior knowledge of the unobserved features, where $T$ is the decision horizon.
@@
The key idea of our proposed algorithm can be summarized as the following two procedures: (i) reconstructing feature vectors via feature augmentation to capture the influence of unobserved features on rewards, and (ii) introducing a novel doubly robust estimator to mitigate information loss due to unobservability. For (i), we decompose the reward into two additive components: one projected onto the row space spanned by the observed features, and the other onto its orthogonal complement. The former term maximally captures the effect of observed features, while the latter minimizes the impact of unobserved features. We then augment observed features with an orthonormal basis from the complementary space, which is orthogonal to the row space of observed features. This formulation enables us to reformulate the problem within a conventional linear bandit framework, where the reward function is defined as a dot product of the augmented features and an unknown parameter, without any additional additive term.
@@
However, since these augmented features are not identical to the unobserved features, potential estimation error may arise due to information loss. To mitigate such errors, we leverage (ii) a doubly robust estimator, which is widely used in the statistical literature for its robustness to errors caused by missing data. These two approaches allow our algorithm to compensate for missing information due to partial observability, improving both estimation accuracy and adaptability to the environment.
@@
  • We propose a linear bandit problem with partially observable features. Our problem setting is more general and challenging than those in the existing literature on linear bandits with latent features, which often rely on specific structural assumptions governing the relationship between observed and latent features. In contrast, our approach assumes no additional structure for the unobserved features beyond the linearity of the reward function, which is commonly adopted in the linear bandit literature (Section 3).
  • We introduce a novel estimation strategy by (i) augmenting the features that maximally captures the effect of reward projected onto the observed features, while minimizing the impact of unobserved features (Section 4), and (ii) constructing a doubly robust (DR) estimator that mitigates errors from unobserved features. By integrating augmented features with the DR estimator, we guarantee a $t^{-1/2}$ convergence rate on the rewards for all arms in each round $t$ (Theorem 2).
@@
  • Our numerical experiments confirm that our algorithm consistently outperforms OLB (Li et al., 2010; Agrawal & Goyal, 2013; Kim & Paik, 2019) and MAB (Lattimore & Szepesvári, 2020) algorithms, validating both its practicality and theoretical guarantees (Section 6).
@@ 2 Related Works
First, our problem setting is more general and challenging than misspecified linear bandit problems, where the reward model deviates from the true reward due to non-linearity (Lattimore & Szepesvári, 2020) or additive deviation terms (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While prior works incorporate cumulative misspecification error into the regret bound, we obtain a regret bound that remains unaffected by such deviations. In particular, Ghosh et al. (2017) employ hypothesis testing for model selection and obtain regret bound of $O(K\sqrt{T}\log T)$ under high misspecification. In contrast, our algorithm achieves $O(\sqrt{(d+d_{h})T\log T})$ regret without such tests, while handling partial observability in a unified framework.
Second, in bandit problems with partially observable features, prior works often rely on structural assumptions. For example, Park & Faradonbeh (2022, 2024); Kim et al. (2023a) assume the true features are drawn from a Gaussian distribution, while Zeng et al. (2025) model them as evolving according to a linear dynamical system with additive Gaussian noise, where observed features are generated via a known linear mapping, also corrupted by Gaussian noise. These methods typically aim to recover the true features—via decoders (Park & Faradonbeh, 2022, 2024), Bayesian oracles (Kim et al., 2023a), or Kalman filtering (Zeng et al., 2025). In contrast, our approach makes no structural assumptions about the relationship between observed and unobserved features and does not attempt to recover the latter. Instead, we mitigate the information loss from partial observability by projecting the latent part of the reward using only the observed features. A more detailed and comprehensive literature review is provided in Appendix A.
@@ 3.1 Notation
For any $n\in\mathbb{N}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. For a vector $\mathbf{v}$, we denote its $L_{1}$, $L_{2}$ and supremum norm by $\|\mathbf{v}\|_{1}$, $\|\mathbf{v}\|_{2}$, and $\|\mathbf{v}\|_{\infty}$, respectively. The $L_{2}$-norm weighted by a positive definite matrix $\mathbf{D}$ is denoted by $\|\mathbf{v}\|_{\mathbf{D}}$. For two vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, the inner product is defined as the dot product, i.e., $\langle{\mathbf{v}_{1}},{\mathbf{v}_{2}}\rangle:=\mathbf{v}_{1}^{\top}\mathbf{v}_{2}$, and we use both notations interchangeably. For a matrix $\mathbf{M}$, its minimum and maximum eigenvalue are denoted by $\lambda_{\min}(\mathbf{M})$ and $\lambda_{\max}(\mathbf{M})$, respectively. We let $\mathrm{R}(\mathbf{M})$ denote the row space of $\mathbf{M}$, i.e., a subspace spanned by the rows of $\mathbf{M}$.
@@ 3.2 Problem Formulation
In this section, we outline our problem setting and introduce several key assumptions. Each arm $a\in[K]$ is associated with a true feature vector $\mathbf{z}_{a}\in\mathbb{R}^{d_{z}}$ that determines the rewards. However, the agent can observe only a subset of its elements, with the others remaining unobserved. Specifically, $\mathbf{z}_{a}$ is defined as follows:
@@
For clarity, we highlight the observed components in blue and the unobserved components in red. Note that the dimensions of the latent feature vector, $d_{u}=d_{z}-d$, and the true feature vector, $d_{z}$, are both unknown to the agent. As a result, the agent is unaware of how many features—if any—are unobserved, or even whether partial observability exists at all. This lack of structural information presents a fundamental challenge, as it prevents the agent from explicitly modeling or compensating for the unobserved components when selecting an appropriate learning strategy.
It is worth noting that the setting with fixed observed features [^2] This assumption is standard in linear bandits with model misspecification (Ghosh et al., 2017; Lattimore et al., 2020), which is a special case of our partially observable feature setting. includes linear bandits with misspecification error (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022) as special cases. In Appendix D, we present a setting with varying observed features and an algorithm that achieves $\sqrt{T}$-rate regret bound. Moreover, if latent features were allowed to change arbitrarily over time, the problem would become non-learnable and thus ill-posed. Consequently, assuming fixed features is both natural and well-justified (see Table 1 for comparisons).
Table 1: Summary of problem settings covered in this paper and the corresponding results. Note that if latent features arbitrarily change over time, the problem itself would become non-learnable, making the problem ill-posed (see Appendix D for details).
Observed Features Unobserved Features Learnable? Results
Fixed Fixed Yes Theorem 3
Varying Fixed Yes Theorem 6
@@
The reward associated with each arm is defined as the dot product of its true features $\mathbf{z}_{a}$ and an unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$, given by
$y_{a,t}=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle+\epsilon_{t}\qquad\forall a\in[K].$ (2)
Here, $\epsilon_{t}$, represents a random noise term that captures the inherent stochasticity in the reward generation process. We provide an assumption on $\epsilon_{t}$, following the assumption commonly adopted in bandit problems.
@@
Let $\{\mathcal{F}_{t}\}_{t\in[T]}$ denote history at round $t$, represented by a filtration of sigma algebras. The reward noise $\epsilon_{t}$ is assumed to be a $\sigma$-sub-Gaussian random variable conditioned on $\mathcal{F}_{t}$. Formally, for all $\lambda\in\mathbb{R}$,
@@
It follows that $\mathbb{E}[{\epsilon_{t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=0$, and $\mathbb{E}[{y_{a,t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ under this assumption. For brevity, we use $\mathbb{E}_{t-1}[\cdot]$ to denote $\mathbb{E}[{\cdot}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]$ henceforth. Given that $\epsilon_{t}$ is sampled after each action is observed, $\epsilon_{t}$ is $\mathcal{F}_{t}$-measurable. To eliminate issues of scale for analysis, we assume that the expected reward $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$ for all $a\in[K]$, and we do not assume any bound on the norm of $\mathbf{z}_{a}$ for all $a\in[K]$.
Let $a_{\star}:=\mathop{\mathrm{argmax}}_{a\in[K]}\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ denote the optimal action, considering both observed and latent features, and $a_{t}$ denote an action chosen in round $t$. The theoretical performance of our algorithm is evaluated through cumulative regret, which measures the total expected difference between the reward of the optimal action and the reward of the chosen action:
$\mathrm{Reg}(T)=\mathbb{E}\left[\sum_{t=1}^{T}\langle{{\mathbf{z}_{a_{\star}}}-{\mathbf{z}_{a_{t}}}},{\boldsymbol{\theta}_{\star}}\rangle\right].$
Considering the composition of $\mathbf{z}_{a}$ defined in Eq. 1, we can decompose the parameter as $\boldsymbol{\theta}_{\star}=[({\boldsymbol{\theta}_{\star}^{(o)}})^{\top},({\boldsymbol{\theta}_{\star}^{(u)}})^{\top}]^{\top}$, where $\boldsymbol{\theta}_{\star}^{(o)}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}\in\mathbb{R}^{d_{u}}$ are the parameters for observed and latent features, respectively. Adopting this decomposition of $\boldsymbol{\theta}_{\star}$, the reward $y_{a_{t},t}$ defined in Eq. 2 can be decomposed into the following three terms:
@@
The last term of Eq. 3 corresponds to the inaccessible portion of the reward. This reward model is equivalent to that imposed in the linear bandits with misspecification error (Lattimore et al., 2020). While the regret bound in Lattimore et al. (2020) includes misspecification error that grows linearly in decision horizon, our proposed method (Section 4) addresses this misspecification error and achieves a regret bound that is sublinear in the decision horizon.
Before presenting our method and algorithm, we first establish a lower bound for the regret incurred by algorithms that disregard the unobserved portion of rewards. Specifically, the following theorem establishes a lower bound for two OLB algorithms: OFUL (Abbasi-yadkori et al., 2011) and LinTS (Agrawal & Goyal, 2013).
@@
Consider a linear bandit problem with an action set $\mathcal{A}:=\{1,2\}$, where the true feature set is defined as $\mathcal{Z}:=\{[1,3]^{\top},[2,19/4]^{\top}\}\subset\mathbb{R}^{2}$ and arm 2 is the optimal action. The latent portion of each reward is “large”, in the sense that there exists a constant $C>0$ such that $|\langle{\mathbf{u}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(u)}}\rangle|=|\mathbb{E}_{t-1}[y_{a_{t},t}]-\langle{\mathbf{x}_{a_{t}}},{\boldsymbol{\theta}_{\star}^{(o)}}\rangle|>C$, following the decomposition described in Eq. 3. Under this setup, the estimator based solely on observed features is not consistent, and the OFUL and LinTS select the suboptimal arm with probability $\Theta(1)$. Consequently, the regret grows linearly in $T$ in both cases. ∎
Theorem 1 implies that neglecting the latent portion of the reward in decision-making could result in a failure in the learning process of the agent. The comprehensive proof is deferred to Section E.1. While Theorem 1 focuses on OFUL and LinTS, which are known to achieve the most efficient regret bounds for UCB and Thompson Sampling-based policies within the linear bandit framework, we prove an algorithm-agnostic lower bound using different analysis (see Appendix F for details).

### v3

@@ Abstract
We study the linear bandit problem that accounts for partially observable features. Without proper handling, unobserved features can lead to linear regret in the decision horizon $T$, as their influence on rewards is unknown. To tackle this challenge, we propose a novel theoretical framework and an algorithm with sublinear regret guarantees. The core of our algorithm consists of (i) feature augmentation, by appending basis vectors that are orthogonal to the row space of the observed features; and (ii) the introduction of a doubly robust estimator. Our approach achieves a regret bound of $\widetilde{O}(\sqrt{(d+d_{h})T})$, where $d$ is the dimension of the observed features and $d_{h}$ depends on the extent to which the unobserved feature space is contained in the observed one, thereby capturing the intrinsic difficulty of the problem. Notably, our algorithm requires no prior knowledge of the unobserved feature space, which may expand as more features become hidden. Numerical experiments confirm that our algorithm outperforms both non-contextual multi-armed bandits and linear bandit algorithms depending solely on observed features.
@@ 1 Introduction
We consider a linear bandit problem where the learning agent has access to only a subset of the features, while the reward is determined using the complete set of features, including both observed and unobserved elements. Conventional linear bandit problems rely on the assumption that the rewards are linear in only observed features, without accounting for the potential presence of unobserved features. However, in many real-world applications, rewards are often affected by the unobserved—hence latent—features that are not observable by the agent. For example, in online advertising without personalization—i.e., when serving general users—each advertisement is associated not only with observable features such as its category, format, or display time, but also with unobservable factors such as emotional appeal, creative design quality, and brand familiarity. These latent factors influence users’ click-through rates, yet they are not directly quantifiable. Similarly, in clinical trials, treatment outcomes depend not only on observable features like dosage and formulation, but also on unobserved factors such as manufacturing variability and potential side effect risks. Hence, accurately incorporating latent features is essential for providing precise recommendations.
@@
To address latent features,  Park & Faradonbeh (2022, 2024); Kim et al. (2023a); Zeng et al. (2025) typically assume that the true features follow a specific distribution, such as a Gaussian distribution. Establishing a sublinear regret bound in the decision horizon without such structural assumptions on the latent features remains a significant challenge and has not been accomplished yet. Key challenges in the bandit problem with partially observable features arise from the complete lack of information about the latent features—indeed, the agent does not even know whether the features are partially observed. We tackle these challenges by proposing a novel linear bandit algorithm that is agnostic to partial observability. Despite having no knowledge of the unobserved features, our algorithm achieves a tighter regret bound than both linear bandit algorithms that consider only observed features and multi-armed bandit (MAB) algorithms that ignore features entirely. For brevity, we will refer to linear bandit algorithms relying solely on observed features as “OLB algorithms” henceforth. Specifically, our proposed algorithm achieves a $\sqrt{T}$-rate regret bound, without requiring any prior knowledge of the unobserved features, where $T$ is the decision horizon.
@@
The key idea of our proposed algorithm can be summarized in the following two procedures: (i) reconstructing feature vectors via feature augmentation to capture the influence of unobserved features on rewards, and (ii) introducing a novel doubly robust (DR) estimator to mitigate information loss due to partial observability. For (i), we decompose the reward into two additive components: one projected onto the row space spanned by the observed features, and the other projected onto its orthogonal complement. The former term maximally captures the effect of observed features, while the latter minimizes the impact of unobserved features. We then augment observed features with an orthonormal basis from the complementary space, which is orthogonal to the row space of observed features. This formulation enables us to reformulate the problem within a conventional linear bandit framework, where the reward function is defined as a dot product of the augmented features and an unknown parameter, without any additional additive term.
@@
However, since these augmented features are not identical to the unobserved features, estimation errors may arise from information loss. To mitigate such errors, we leverage (ii) a DR estimator, which is widely used in the statistical literature for its robustness to errors caused by missing data. These two approaches allow our algorithm to compensate for missing information due to partial observability, improving both estimation accuracy and adaptability to the environment.
@@
  • We propose a linear bandit problem with partially observable features. Our problem setting is more general and challenging than those in the existing literature on linear bandits with latent features, which often rely on specific structural assumptions governing the relationship between observed and latent features. In contrast, our approach assumes no additional structure for the unobserved features beyond the linearity of the reward function, which is commonly adopted in the linear bandit literature (Section˜3).
  • We introduce a novel estimation strategy by (i) augmenting the features that maximally capture the effect of reward projected onto the observed features, while minimizing the impact of unobserved features (Section˜4), and (ii) constructing a DR estimator that mitigates errors from unobserved features. By integrating augmented features with the DR estimator, we guarantee a $\sqrt{t}$-convergence rate on the rewards for all arms in each round $t$ (Theorem˜2).
@@
  • Our numerical experiments confirm that our algorithm consistently outperforms OLB (Li et al., 2010; Agrawal & Goyal, 2013; Kim & Paik, 2019) and MAB (Lattimore & Szepesvári, 2020) algorithms, validating both its practicality and theoretical guarantees (Section˜6).
@@ 2 Related Works
First, our problem setting is more general and challenging than misspecified linear bandit problems, where the reward model deviates from the true reward due to non-linearity (Lattimore & Szepesvári, 2020) or additive deviation terms (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022). While prior works incorporate cumulative misspecification error into the regret bound, we obtain a regret bound that remains unaffected by such deviations. In particular, Ghosh et al. (2017) employ hypothesis testing for model selection and obtain a regret bound of $O(K\sqrt{T}\log T)$ under high misspecification. In contrast, our algorithm achieves a regret bound of $O(\sqrt{(d+d_{h})T\log T})$ without such tests, while handling partial observability in a unified framework.
Second, in bandit problems with partially observable features, prior works often rely on structural assumptions. For example, Park & Faradonbeh (2022, 2024); Kim et al. (2023a) assume the true features are drawn from a Gaussian distribution, while Zeng et al. (2025) model them as evolving according to a linear dynamical system with additive Gaussian noise, where observed features are generated via a known linear mapping, also corrupted by Gaussian noise. These methods typically aim to recover the true features—via decoders (Park & Faradonbeh, 2022, 2024), Bayesian oracles (Kim et al., 2023a), or Kalman filtering (Zeng et al., 2025). In contrast, our approach makes no structural assumptions about the relationship between observed and unobserved features and does not attempt to recover the latter. Instead, we mitigate the information loss from partial observability by projecting the latent part of the reward using only the observed features. A more detailed and comprehensive literature review is provided in Appendix˜A.
@@ 3.1 Notation
For any $n\in\mathbb{N}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. For a vector $\mathbf{v}$, we denote its $L_{1}$, $L_{2}$ and supremum norms by $\|\mathbf{v}\|_{1}$, $\|\mathbf{v}\|_{2}$, and $\|\mathbf{v}\|_{\infty}$, respectively. The $L_{2}$-norm weighted by a positive definite matrix $\mathbf{D}$ is denoted by $\|\mathbf{v}\|_{\mathbf{D}}$. For two vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$, the inner product is defined as the dot product, i.e., $\langle{\mathbf{v}_{1}},{\mathbf{v}_{2}}\rangle:=\mathbf{v}_{1}^{\top}\mathbf{v}_{2}$, and we use both notations interchangeably. For a matrix $\mathbf{M}$, its minimum and maximum eigenvalues are denoted by $\lambda_{\min}(\mathbf{M})$ and $\lambda_{\max}(\mathbf{M})$, respectively. We let $\mathrm{R}(\mathbf{M})$ denote the row space of $\mathbf{M}$, i.e., the subspace spanned by the rows of $\mathbf{M}$.
@@ 3.2 Problem Formulation
In this section, we outline our problem setting and introduce several key assumptions. Each arm $a\in[K]$ is associated with a true feature vector $\mathbf{z}_{a}\in\mathbb{R}^{d_{z}}$ that determines its rewards. However, the agent can observe only a subset of its elements, with the others remaining unobserved. Specifically, $\mathbf{z}_{a}$ is defined as follows:
@@
For clarity, we highlight the observed components in blue and the unobserved components in red. Note that the dimensions of the latent feature vector, $d_{u}=d_{z}-d$, and the true feature vector, $d_{z}$, are both unknown to the agent. As a result, the agent is unaware of how many features, if any, are unobserved, or even whether partial observability exists at all. This lack of structural information presents a fundamental challenge, as it prevents the agent from explicitly modeling or compensating for the unobserved components when selecting an appropriate learning strategy.
It is worth noting that the setting with fixed observed features [^2] This assumption is standard in linear bandits with model misspecification (Ghosh et al., 2017; Lattimore et al., 2020), which is a special case of our partially observable feature setting. includes linear bandits with misspecification error (Ghosh et al., 2017; Bogunovic et al., 2021; He et al., 2022) as special cases. In Appendix˜D, we present a setting with varying observed features and an algorithm that achieves a $\sqrt{T}$-rate regret bound. Moreover, if latent features were allowed to change arbitrarily over time, the problem would become non-learnable and thus ill-posed. Consequently, assuming fixed features is both natural and well-justified (see Table˜1 for comparisons).
The reward associated with arm $a$ is defined as the dot product between its true feature vector $\mathbf{z}_{a}$ and an unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{d_{z}}$, namely
$y_{a,t}=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle+\epsilon_{t}\qquad\forall\ a\in[K],$ (2)
where $a_{t}\in[K]$ denotes the action selected by the agent at round $t$. Here, $\epsilon_{t}$ represents a random noise term that captures the inherent stochasticity in the reward generation process. We make the following assumption on $\epsilon_{t}$, which is commonly adopted in bandit problems.
Table 1: Summary of problem settings covered in this paper and the corresponding results. Note that if latent features arbitrarily change over time, the problem itself would become non-learnable, making the problem ill-posed (see Appendix˜D for details).
Observed
Features
Unobserved
Features
Learnable? Results
Fixed Fixed Yes Theorem˜3
Varying Fixed Yes Theorem˜6
@@
Let $\mathcal{F}_{t}$ denote the history at round $t$, represented by a filtration of $\sigma$-algebras, e.g., $\sigma(a_{1},y_{a_{1},1},\ldots,a_{t-1},y_{a_{t-1},t-1},a_{t})$. The reward noise $\epsilon_{t}$ is conditionally $\sigma$-sub-Gaussian, i.e., for all $\lambda\in\mathbb{R}$,
@@
Since $\epsilon_{t}$ is sampled after $a_{t}$ is observed, $\epsilon_{t}$ is $\mathcal{F}_{t}$-measurable. Under Assumption˜1, it follows that $\mathbb{E}[{\epsilon_{t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=0$, thus $\mathbb{E}[{y_{a,t}}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]=\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$. For brevity, we use $\mathbb{E}_{t-1}[\cdot]$ to denote $\mathbb{E}[{\cdot}{\hskip 0.80002pt|\hskip 0.80002pt}{\mathcal{F}_{t-1}}]$ henceforth. To eliminate issues of scale for analysis, we assume that the expected reward $|\langle\mathbf{z}_{a},\boldsymbol{\theta}_{\star}\rangle|\leq 1$ for all $a\in[K]$, and we do not assume any bound on the norm of $\mathbf{z}_{a}$ as well.
Let $a_{\star}:=\mathop{\mathrm{argmax}}_{a\in[K]}\langle{\mathbf{z}_{a}},{\boldsymbol{\theta}_{\star}}\rangle$ denote the optimal action, considering both observed and latent features. Moreover, we denote by $\mathbf{z}_{\star}$ and $y_{\star,t}$ the true feature vector and the realized reward associated with the optimal action $a_{\star}$, respectively. We evaluate the theoretical performance of our algorithm using cumulative regret, which is defined as the expected sum of the differences between the reward of $a_{\star}$ and $a_{t}$:
$\mathrm{Reg}(T)$ $=\sum_{t=1}^{T}\langle{\boldsymbol{\theta}_{\star}},{\mathbf{z}_{\star}-\mathbf{z}_{a_{t}}}\rangle$
$=\sum_{t=1}^{T}\left(\mathbb{E}_{t-1}[y_{\star,t}]-\mathbb{E}_{t-1}[y_{a_{t},t}]\right).$
Considering the composition of $\mathbf{z}_{a}$ defined in Eq.˜1, we decompose the parameter vector as $\boldsymbol{\theta}_{\star}=[({\boldsymbol{\theta}_{\star}^{(o)}})^{\top},({\boldsymbol{\theta}_{\star}^{(u)}})^{\top}]^{\top}$, where $\boldsymbol{\theta}_{\star}^{(o)}\in\mathbb{R}^{d}$ and $\boldsymbol{\theta}_{\star}^{(u)}\in\mathbb{R}^{d_{u}}$ correspond to the parameters associated with the observed and latent features, respectively. Adopting this decomposition of $\boldsymbol{\theta}_{\star}$, the reward $y_{a_{t},t}$ defined in Eq.˜2 can be decomposed into the following three terms:
@@
The last term of Eq.˜3 corresponds to the inaccessible portion of the reward. This reward model is equivalent to that imposed in the linear bandits with misspecification error (Ghosh et al., 2017; Lattimore et al., 2020). While the regret bound in Lattimore et al. (2020) includes misspecification error that grows linearly with decision horizon, our proposed method (Section˜4) addresses this misspecification error and achieves a regret bound that is sublinear in $T$.
Before presenting our method and algorithm, we first establish a lower bound on the regret of linear bandit algorithms that disregard the unobserved portion of the rewards. In particular, the following theorem provides lower bounds for two OLB algorithms: OFUL (Abbasi-yadkori et al., 2011) and LinTS (Agrawal & Goyal, 2013).
@@
Consider a linear bandit problem instance with an action set $\mathcal{A}:=\{1,2\}$, where $d=d_{u}=1$, implying that $d_{z}=2$. The true feature vectors are given by $\mathcal{Z}:=\{[1,3]^{\top},[2,19/4]^{\top}\}\subset\mathbb{R}^{2}$, with the agent observing only the first element of each vector, while the second element remains unobserved. In this setting, we take arm 2 to be the optimal action. However, an estimator relying solely on the observed features is inconsistent, leading OFUL and LinTS to select the suboptimal arm with probability $\Theta(1)$, and thereby incur regret that grows linearly in $T$. ∎
Theorem˜1 shows that neglecting the latent portion of the reward in decision-making may cause the agent to fail in identifying the optimal action. The comprehensive proof is deferred to Section˜E.1. While Theorem˜1 specifically considers OFUL and LinTS—which are known to achieve the most efficient regret bounds for UCB- and Thompson sampling-based policies in the linear bandit framework—we additionally present an algorithm-agnostic lower bound based on a different analysis (see Appendix˜F for details).